OpenStudy (anonymous):

the function below crosses the x-axis twice true or false 2x^2-3x+1

OpenStudy (debbieg):

Under what condition will a quadratic cross the x-axis twice? What do you know about the roots, if it crosses twice?

OpenStudy (anonymous):

its positive huh

OpenStudy (debbieg):

What's positive?

OpenStudy (debbieg):

(You might be right - I just need to see exactly what you mean by that.)

OpenStudy (anonymous):

if it crosses twice its positive

OpenStudy (debbieg):

OK, "its" is very vague here, lol. I asked you: what does the fact that a quadratic crosses the x-axis twice tell you about the roots of the quadratic?

OpenStudy (debbieg):

There are 3 possibilities for the roots of a quadratic: 2 real roots 1 repeated real root (something like \((x-3)^2=0\) means 3 is a repeated root or no real roots (2 imaginary roots) Each of those situations has a certain "look" to the graph, in terms of where/if/how many times the function crosses the x-axis. I'm asking you, if you understand that connection.

OpenStudy (debbieg):

And there is SOMETHING that tells you WHICH of those 3 situations you are in. Do you know what that something is?

OpenStudy (anonymous):

ohhhhh its gonna be false cuz theres a negative

OpenStudy (debbieg):

I'm still not sure what you mean. What's a negative??

OpenStudy (anonymous):

becuse if it negative it cross the x-axis zero times

OpenStudy (debbieg):

Remember the discussion yesterday about crossing/touching and the discriminant? This is closely related to that.

OpenStudy (debbieg):

IF WHAT'S NEGATIVE?

OpenStudy (debbieg):

http://openstudy.com/study#/updates/52153cdae4b0450ed75e459c If discr>0, there are 2 DISTINCT real roots - what does that mean in terms of it's behavior crossing/touching the axis? If discr=0, there is ONE REPEATED real root - what does that mean in terms of it's behavior crossing/touching the axis? if discr<0, there are NO REAL roots (only complex roots) - what does that mean in terms of it's behavior crossing/touching the axis?

OpenStudy (debbieg):

@romanortiz65 what's the value of the discriminant for this function?

OpenStudy (anonymous):

-5 or +1

OpenStudy (debbieg):

There can only be on discriminant. How are you getting 2 values? Show me your work.

OpenStudy (debbieg):

*one

OpenStudy (anonymous):

1

OpenStudy (debbieg):

Right! So how many real roots are there?

OpenStudy (anonymous):

two bu what i think the formula is b^2-4ac

OpenStudy (debbieg):

Exactly - discrim=(-3)^2-4(2)(1)=9-8=1>0 That means 2 real roots. And that means it crosses the x-axis how many times?